### v1

# VisFly: An Efficient and Versatile Simulator for Training Vision-based Flight

Abstract
State-of-art simulators primarily focus on providing full-stack simulation tools or state-only parallelizability. Due to the limitation of computing resources, they have to make trade-off among photo-realism and sampling efficiency. Yet, both factors are crucial for data-driven reinforcement learning tasks. Therefore, we introduce a both rapid-rendering and photo-realistic quadrotor simulator: VisFly. VisFly offers a user-friendly framework and interfaces for users to develop or utilize. It couples differentiable dynamics and habitat-sim rendering engines, reaching frame rate of up to 10000 frame per second in cluttered environments. The simulation is wrapped as a gym environment, facilitating convenient implementation of various baseline learning algorithms. It can directly import all the open-source scene datasets compatible with habitat-sim, which provides more fair benchmarks for comparing the intelligent policy. VisFly presents a general policy architecture for tasks, and the whole framework is verified by three regular quadrotor tasks with visual observation. We will make this tool available at https://github.com/SJTU-ViSYS/VisFly.
Index Terms:
Quadrotor Simulation, Vision-based Learning, Gym-wrapped Framework, Benchmark Visual Tasks.
I Introduction
Quadrotors have great potential applications in various tasks such as search-and-rescue, surveillance, and inspection operations. The capability of autonomous flight of quadrotors is the key to enabling them to accomplish complex tasks, especially in challenging environments where GPS signals are unavailable and full of obstacles. Typically, an autonomous flight stack consists of localization, environmental perception, planning, and control components. Significant progress has been made in each component that allows quadrotors being able to autonomously navigate in cluttered environments [1, 2] with only onboard sensory and computational resources. However, the classic pipeline heavily relies on the performance of individual components. The accumulation of errors and running time of individual components could greatly reduce the agility and robustness of autonomous flight [3].
An alternative approach is to train a neural network (policy) to receive the sensory input and directly send the control commands to the quadrotor. Such methods, also known as the learning-based approach, typically integrate several modules into a single flight model, offering the potential to achieve autonomy with low latency and high robustness. One way to train the flight policy is to collect real-world data by human-pilot drones [4, 5, 6], but this approach is inefficient and prohibitively expensive. Alternatively, simulation can generate datasets, and policies trained on these simulated datasets often achieve performance comparable to those trained on real-world data[7]. Researchers [3] have demonstrated impressive collision avoidance capabilities by employing an algorithmic expert to gather training data for the flight policy. However, designing such experts to achieve both algorithmic optimality and efficiency can be challenging, thereby limiting the applicability of this approach to complex tasks.
Recently, researchers have demonstrated that reinforcement learning (RL) facilitates not only high-level decision-making applications such as autonomous driving [8], strategy games[9], but also basic control problems for complicated systems, like humanoid [10], quadruped locomotion [locomotion]. These potential and powerful applications inspire academics in the field of aerial robotics to explore reinforcement learning to enhance the autonomous flight [11, 12]. RL-based approach usually relies a simulator to collecting massive data generated from the interaction between the agent and the simulated environment. There are already well-designed simulators [13, 14] widely used in the robotics community, with some specifically dedicated to quadrotors [15, 16]. These simulators are primarily designed for fast prototyping and iteration, typically encompassing the full stack of hardware and software simulations with an emphasis on closed-loop validation and testing. They usually operate at slow frame rates (10 to 100 fps), which does not negatively impact testing or validation but presents significant drawbacks for reinforcement learning due to low sampling efficiency. Some quadrotor simulators [17, 18, 19] have addressed these drawbacks by improving the sampling of physical state data, enabling the training of state-based policies for autonomous flight. However, they still lack the capability to sample vision inputs at high frame rates, making it challenging to train a high-performance vision-based flight policy.
Figure 1: Overall diagram of VisFly. The differentiable physics and Habitat-sim rendering engine constitute the quadrotor agents, which is then wrapped as gym environments, preserving interfaces for RL algorithms.
Given that vision is one of the most informative and cost-effective sources for perception, it holds substantial potential for deploying a vision-based autonomous flight model in real-world applications. The key to effectively training a vision-based flight policy for quadrotors via reinforcement learning lies in 1) the rendering speed of the vision data, 2) efficient computation of flight physics, 3) the diversity of scenes for training, and 4) the adoption of state-of-the-art reinforcement learning algorithms. Moreover, to enable the autonomous flight of a team of quadrotors, multi-agent learning should also be supported.
To achieve these goals, we developed an efficient simulation environment for training vision-based autonomous flight of quadrotors. Our environment leverages the PyTorch, Habitat-Sim [20], and OpenAI’s Gym [21] platforms. We implemented flight dynamics and three widely-used flight controllers using PyTorch, making flight physics not only computationally efficient but also differentiable, thereby enabling the use of model-based algorithms with known dynamics [22]. By adopting Habitat-Sim’s rendering engine, our environment can achieve high frame rates at 10,000 frames per second while rendering color images at a resolution of 64×64, which is about two orders of magnitude faster than existing quadrotor simulators. Moreover, our simulator can load a series of open-source datasets comprising a vast number of 3D scene models (including real-world 3D scans), allowing us to train vision-based flight policies in diverse and realistic scenes. Our simulator is highly flexible, featuring Gym-wrapped interfaces for integrating the latest learning algorithms and customized tasks. With our well-designed parallel mechanism, we can run multiple agents within different scenes simultaneously, facilitating training vision-based flight policies in multi-agent settings.
To validate our environment, we present an example of using a general network architecture to train vision-based flight policies for three autonomous flight tasks: landing, navigation, and cooperative flight through narrow gaps. All tasks rely solely on vision sensory inputs without any external guidance. The results demonstrate that our simulation environment can be used to train vision-based flight policies in both single and multi-agent settings with high success rates in a very short period. Our simulation environment will be released as an open-source project and will be continuously updated. To summarize, the contributions of this work include:
  1. 1.
    We present a novel quadrotor simulator that achieves high performance of rendering both physics and vision data, allowing to train vision-based flight policies efficiently.
  2. 2.
    Our simulator supports single-agent and swarm scenarios in parallelized physics and scenes, with Gym-wrapped interfaces for implementing the latest learning algorithms and customization of tasks. It also integrates open-source datasets for diverse scenes.
  3. 3.
    We demonstrate the effectiveness of our training environment for training diverse vision-based tasks, highlighting its versatility.
Table I: Key Character Comparison of Popular Simulators
Engine Parallel FPS Sensor Gym- 3D Differ- Swarm
Physics Render agent scene Physics Render wrap Datasets entiable
RotorS Gazebo OpenGL $1\times 10^{2}$ 20 I, D, C
Airsim PhysX Unreal $1\times 10^{2}$ 60 I, D, C, S
Flightmare Flexible Unity $2.2\times 10^{5}$ $2.3\times 10^{2}$ I, D, C, S
PyBulletDrone Bullet OpenGL $1\times 10^{4}$ $3\times 10^{3}$ I, D, C, S
FlightGoggles Flexible Unity $1\times 10^{3}$ 64 I, D, C, S
CrazyS Gazebo OpenGL $1\times 10^{3}$ 20 I, D, C
OmniDrone Self-defined Isaac Sim $2\times 10^{5}$ I, D, C, S
FastSim Flexible Unity $1\times 10^{2}$ 40 I, D, C, S
Ours Self-defined Habitat-sim $3.7\times 10^{4}$ $1\times 10^{4}$ I, D, C, S
  • 1
    I, C, D, S refer to IMU, RGB, depth, segmentation. The simulators with superscript * are full-stack, while the others are learning-specialized. VisFly is light-weight, which does not have too much requirements for hardware. Noting that, FPS data provided by original researches are not evaluated with same number of agents, but it still reflect the performance to a large extent. ”-” in FPS column means this function is unavailable. All the frame rates are test with 100 environments if parallelization is available. Render FPS are test while rendering $64\times 64$ depth images. If the simulator cannot be deployed locally at present, FPS will be directly cited from original paper.
II Related Work
II-A Full-stack simulators
The initial aim of simulators is to create a simulated platform for users to verify the algorithmic prototypes because quadrotors with immature algorithm are easy to crash. This category always consists of comprehensive full-stack simulated flight modules such as perception, planning, and control, especially for highly realistic physics. Gazebo [13] is a widely used full-stack robotics simulator of this category. While it is general-purpose simulator, it has strong support for UAVs through various plugins and models. Considering various requirements, researchers develop simulators specific for UAVs based on Gazebo. RotorS [15] is a popular multi-rotor UAV modular simulator built on Gazebo, provides flexible interfaces to replace several parts of the pipeline, making it convenient to verify state estimators or control algorithms. Rather than using open-source rendering engines as in Gazebo, several simulators utilize commercial rendering engines such as Unity  [^1] https://unity.com/ and Unreal Engine  [^2] https://www.unrealengine.com/ to achieve high visual fidelity. For instance, FlightGoogles[23] employs the Unity engine for realistically simulating image sensors and uses real-world flight data to create avatars in simulation environments, enabling vehicle-in-the-loop simulation. Similarly, AirSim [14] is built on Unreal Engine to offer physically and visually realistic environments, along with real-time hardware-in-the-loop (HITL) simulation capabilities. Additionally, AvoidBench [24], built on the Unity engine, offers valuable benchmarks to evaluate obstacle avoidance algorithms. FastSim [25], also built on Unity, stands out as the first to integrate path planning algorithms directly into the simulator, facilitating tests of high-level autonomous control functionalities. While using commercial engines significantly enhances image quality (including depth maps and semantic segmentation), it comes at the cost of slower rendering speed because of their sophisticated rendering pipelines, particularly for the vision data. However, for applications focusing on learning in simulators using vision sensors, it is critical to acquire vision data quickly for fast training.
II-B Learning-specialized simulators
High acquisition rates for both sensory data and robot physics significantly accelerate the training process of learning-based algorithms. This makes them essential features for simulators designed for learning. Notably, physics computation can be achieved at very high frequencies in some existing simulators. For instance, by paralleled running, the frame rate of physics computation can reach an impressive 100000 FPS in Flightmare [17] and OminiDrone [18]. However, the frame rates will drop down to hundreds of FPS if the visual observation is rendered simultaneously in Flightmare. While PyBulletDrone [19] achieves high rendering FPS, it focuses more on scene visualization rather than sensor data acquisition. Additionally, the basic OpenGL-based rendering engine (TinyRenderer for cpu rendering or OpenGL3 for GPU rendering ) lacks the functionality to import complex scenes and objects, making it difficult for training navigation policies using diverse and complicated scene data.
Selecting a rendering platform for training vision-based flight policies necessitates balancing speed, open-source scene availability, and hardware requirements. Habitat-sim[20], a lightweight simulator for training embodied agents, excels in rendering speed due to its customized rendering engine. It is compatible with numerous popular open-source datasets, including photo-realistic 3D reconstructions and digital forms, and offers flexible interfaces for customizing scenes and agent configurations, all with minimal hardware performance requirements. However, it focuses solely on ground robots and manipulations. In contrast, Nvidia’s Isaac Sim  [^3] https://developer.nvidia.com/isaac/sim is a comprehensive simulator that features extensive joint optimizations in rendering and simulation at the software architecture level. Nevertheless, it has higher hardware requirements and provides less flexibility and extensibility compared to Habitat-sim. Given these considerations, Habitat-sim appears to be a promising tool for addressing the limitations of existing learning-specialized simulators for aerial robots.
Our simulator is built on Habitat-sim, utilizing its fast rendering engine and well-designed interfaces to leverage existing massive 3D datasets. However, unlike Habitat-sim, which focuses on ground robots and manipulations, our simulator is tailored specifically for aerial robots. For the convenience of implementation, existing learning-specialized simulators, such as Flightmare, OminiDrone, PyBulletDrone wrap the environment upon OpenAI’s gym environment standards. An OpenAI’s gym class receives an action and updates its internal state, and returns a new state paired with a corresponding reward. [TODO] but the examples are highly customized for specific tasks. Each of them provides individual environments or scenes, making it hard to evaluate the performance at a common and fair metric. Therefore, we also adopt a modular design for integrating the Gym framework in our simulator to make it easier to customize the vision-based flight tasks. To highlight the novel features of our simulator, we compare it with existing simulators in Table I.
III Methodology
VisFly consists of three parts: (1) Differentiable, and parallelized dynamics. (2) Rapid-rendering, parallelized, photo-realistic scene manager built by habitat-sim. (3) Gym-wrapped interfaces and its extensions. A schematic diagram of VisFly is shown in Figure 1.
Figure 2: RGB, depth, segmentation photos captured in HM3D, MP3D, HSSD, Replica, and customized scenes by generation. The right subfigure is the swarm of quadrotors in a clear garage.
III-A Dynamics
The dynamics simulation is implemented using PyTorch, allowing physics to run on GPUs. Furthermore, PyTorch’s computational graph enables the computation of analytical gradients through backpropagation, which makes it a differentiable process. It could be directly leveraged as a known dynamic transition model in model-based RL [26]. Analytical gradients can descend along much more precise directions because it follows the precise dynamic process, thereby accelerating training several times compared to common value-based or policy-based reinforcement learning [22, 27]. Considering general multi-rotors dynamics, we build the quadrotor model as:
$\dot{\mathbf{x}}_{W}={\mathbf{v}}_{W}$ (1)
$\dot{\mathbf{v}}_{W}=\frac{1}{m}\mathbf{R}_{WB}(\mathbf{f}+\mathbf{d})+\mathbf{g}$
$\dot{\mathbf{q}}=\frac{1}{2}\mathbf{q}\otimes\mathbf{\Omega}$
$\dot{\mathbf{\Omega}}=\mathbf{J}^{-1}(\eta-\mathbf{\Omega}\times\mathbf{J}\mathbf{\Omega})$
where $\mathbf{x}_{W}$, $\mathbf{v}_{W}$, $\mathbf{q}$ denote position, velocity, orientation (quaternion) of quadrotors in world coordinate, and $\mathbf{\Omega}$ denotes angular velocity in body coordinate. $\otimes$ is multiplication between quaternion and vector. $\mathbf{R}_{WB}$ is the matrix of transformation from body coordinate to world coordinate. $\mathbf{f}$, $\mathbf{\eta}$ denote the collective rotor thrust and momentum. $\mathbf{J}$ indicates the diagonal inertia matrix, while $m$ indicates the quadrotor mass. $\mathbf{d}$ represents the air drag in the body coordinate, assumed positive with the square of velocity [28]:
$\mathbf{d}=\frac{1}{2}\rho\mathbf{v}_{B}\odot\mathbf{v}_{B}\mathbf{C}_{d}\odot\mathbf{s}$ (2)
where $\rho$, $\mathbf{C}_{d}$, $\mathbf{s}$, $\mathbf{v}_{B}$ denote air density, drag coefficients, crossing area, velocity in body coordinates. $\odot$ refers to element-wise multiplication. $\mathbf{f}$, $\mathbf{\eta}$ are computed from single rotor thrust $\mathbf{f}_{i}$:
$\mathbf{f}=\sum_{1}^{4}f_{i}\qquad\mathbf{\eta}=\sum_{1}^{4}\mathbf{T}_{i}\times\mathbf{f}_{i}$ (3)
$\mathbf{T}_{i}$ denotes the moment arm matrix, with $i$ representing the rotor index. To account for control input delay, we introduce the time decay constant $c$ and model it as an exponential process:
$\mathbf{f}_{i}(3)=k_{2}\omega_{i}^{2}+k_{1}\omega_{i}+k_{0}\qquad\omega_{i}=\omega_{i}^{des}+(\omega_{i}^{\prime}-\omega_{i}^{des})e^{-ct}$ (4)
where $\omega$ denotes the rotor speed, and superscripts , $des$ separately indicate current and desired state. $k_{2,1,0}$ are constant thrust coefficients w.r.t. rotor speed. Based on [29, 30], we develop three widely-used control methods, individual single-rotor thrusts (SRT), mass-normalized collective thrust and bodyrates (CTBR), position commands and yaw (PS), and linear velocity commands and yaw (LV). SRT and CTBR are end control inputs for Betaflight, which can be used for training efficient end2end policies[31]. The minimal simulation time step can be set several times shorter than the control time step, resulting in more precise physics. Both 4th order Runge-Kutta and Euler methods are supported, providing users the flexibility to choose based on their platform’s performance.
III-B Scenes
We use Habitat-sim as the engine to render the visual observation of AI agents. Rapid rendering is one of the most crucial character of habitat-sim. We evaluate the frame rate of VisFly while running with physics, and show the result in Figure 3. With visual input, VisFly can reach the frame rate of up to 10000 FPS, almost 100 times of those using Unity or Unreal Engine.
Figure 3: Frame rate of VisFly. We evaluate the frame rate using a desktop with RTX 4090 and 32-core 13th Gen Intel(R) Core(TM) i9-13900K. The multiple agents run repeatedly with random actions. Left sub-figure is the result tested in Replica datasets while total number of agents is 100. The maximum frame rate can reach as high as 10,000 FPS. When the resolution is set to 256$\times$256, a popular resolution choice in open-source models and hardware, the frame rate can also reach up to 6,000 FPS. The right sub-figure illustrates the trend of how the frame rate varies with the number of scenes running simultaneously when resolution is 64$\times$64. Increasing number of scenes decreases the simulating speed while the degradation can be tolerated.
Meanwhile, a series of datasets including HM3D[32], MP3D[33], Replica[34], and HSSD[35] are compatible with habitat-sim. In these datasets, some of the data scanned from the real world are highly photo-realistic. Users can directly utilize these open-source datasets to evaluate or train models. Except for open datasets, we also define datasets generator to generate basic cluttered environments for common usage like training navigation policy. These real-world scenarios, combined with powerful scene customizable features, make VisFly more realistic, effectively reducing the adaptation challenges for robots during Sim2Real transitions. Rich extant datasets eliminate the extra efforts for gathering the original data, and offer a fairer approach compared to self-defined environments and promising to serve as baselines for quadrotor tasks. VisFly supports users to equip quadrotors with several different types of sensors, reserving extension for developing complicated multi-camera aerial robots.
Specialized for ground robots, habitat-sim assumes that robots always operate on a 2-dimensional plane. It precomputes the NavMesh, which projects spatial objects onto the floor and generates distance information to the closest obstacle. This approach is not suitable for aerial robots that move freely. Therefore, we have developed a spatial distance computation module for the modified habitat-sim, enabling it to detect the relative position of the nearest collision point. The modified version of habitat-sim will also be released within the same website.
VisFly supports parallelized agents as well as multi-agents in one scene. These two modes share the same underlying framework. The key difference in outcomes is that the former visually omits other drones and ignores collisions between them, whereas the latter does the opposite.
III-C Sim-to-Real Transfer
The gap between simulation and reality hinders the usage transfer of robots trained in simulation. In order to upgrade the performance of practical implementation, one effective approach is to improve the realism of simulation, whose drawback is discussed in Introduction. The other approach is to enlarge the data distribution generated by simulator, strengthening the in-distribution generalization of model, which is called as domain randomization[7]. Considering the real flight scenarios, we add customizable noise to all the sensors as Table II.
Table II: Noise model of sensors
Sensor Noise Model
IMU Normal
RGB Normal, Poisson, Salt&Pepper, Speckle
Depth Normal, Poisson, Salt&Pepper, Speckle, RedWood[36]
Segmentation Normal, Poisson, Salt&Pepper, Speckle
The initial position, velocity, orientation, angular velocity, can also be randomized to follow a normal or uniform distribution. To avoid unstable training, agents are reset with initial position away from obstacles in cluttered environments. An important contribution of VisFly is that multiple scenes are able to run distributively and return observations centrally together. During training, scenes are treated as a datasets, that means, scenes could be sampled and applied with shuffle option. To the best of our knowledge, VisFly is the first simulator that runs multiple scenes simultaneously to achieve scene randomization.
III-D Gym-wrapped Environment
Gym is an open-source library of several regular interactive environments designed for reinforcement learning. It wraps the processes of environment and set interfaces for user to interact with them. Users are able to obtain the agents state including reward, observation, finished status, running information through step() function.
VisFly’s ultimate mission is to facilitate the development of aerial robots learning for several tasks. Therefore, we have taken the extensibility of the basic environment into account as much as possible when packaging the environment. Two abstract environment class droneGymEnvsBase, multiDroneGymEnvsBase are provided. When users design custom tasks for a quadrotor, they need only define a class that inherits from the abstract gym environment class and override three functions: get_success(), get_reward(), and get_observation() as shown in Algorithm 1. This approach creates a new gym environment tailored to the specific mission. We provide three examples in Section IV for users to reference. This customizable architecture enables the training of models that can act as backbones for more complex tasks. For instance, a model trained for obstacle avoidance can serve as the baseline or backbone for higher-level autonomous tasks, such as search and rescue operations in wilderness areas.
Algorithm 1 High-level Tasks Definition
class DroneGymEnvsBase
         @abstractmethod
         def get_reward()
         @abstractmethod
         def get_observation()
         @abstractmethod
         def get_success()
class DroneNavigationEnvs(DroneGymEnvsBase)
         def get_reward()
               return $r_{distance}+r_{speed}-r_{collision}$
         def get_observation()
               return $state+vision+target$
         def get_success()
               return $status_{reach\_the\_target}$
class DroneLandingEnvs(DroneGymEnvsBase)
         def get_reward()
               return $-r_{height}+r_{speed}-r_{collision}$
         def get_observation()
               return $state+vision+visual\_target$
         def get_success()
               return $status_{height<=0.1}\&status_{speed<=0.1}$
IV Applications
In order to verify this simulator, we list three instances: (1) learning to land, (2) learning to navigate in cluttered environment, (3) learning to cooperatively cross the narrow gap. We choose reinforcement learning to train an end2end model since RL is data-hungry. In the following instances, considering that the demonstration point is relative with training efficiency instead of algorithm categories, we only employ PPO [37] from stable-baselines3[38] (SB3) for its outstanding capacity to handle high-dimensional observation, to train the models. The backbone architecture of the policy network is consistent with SB3 as Figure 4, and we modify it to make it more accessible for any tasks.
The image extractor is a conventional neural network that comprises three layers, each containing a convolutional layer, an activation function, a max pooling operation, and concludes with a linear layer with 128 units. The convolutional layers have kernel sizes of 5, 3, and 3, with corresponding channel counts of 6, 12, and 18. Each max pooling operation uses a kernel size of 2. The target extractor, state extractor, swarm extractor are all multilayer perceptrons (MLPs) that comprise two linear layers with unit counts of 128, 64. Both action net and value net are also MLPs that comprise two linear layers with unit counts of 64, 64. All the activation function are ReLU. Noting that, this architecture has been parameterized, thus making it convenient to fine-tune. The action types in all the instances are CTBR. And the quadrotor take its full state (position, orientation, linear velocity, angular velocity), visual image, target as observation. Most of training hyper-parameters are set as default, detailed setting, as well as reward function, done function, observation function, are introduced in source code.
Figure 4: Neural network architecture illustration of policies. The image feature extractor has been modified to be more customizable, including backbones such as ResNet[39], MobileNet[40], EfficientNet[41]. The policies for three tasks follows this architecture but slightly different. VisFly additionally retains the recurrent network interface [42] for training tasks that do not strictly conform to Markov process.
IV-A Learning to Land
Quadrotor begins from aerial flight status, decelerates, and eventually lands stably on the ground. It continuously monitors the limited landing area directly below. This landing area is a black rectangle positioned horizontally on the ground within a clear garage, as shown in Figure 5. Quadrotor calculates the center of the $0.5\times 0.5m^{2}$ black rectangle in their field of view, strives to keep it centered, and incorporates it as part of their observations. During flight, quadrotor only acquires $64\times 64$ RGB image from camera instead of any precise spatial location of landing area. The center of landing area is calculated using threshold segmentation and incorporated as target observation. This training takes 1.7 hour, and the additional training time from the Section IV-B task accounts for the center preprocessing required for the landing area.
Figure 5: Training results of landing, navigation, and swarm crossing. These subfigure columns respectively represent the results of landing, navigation, swarm crossing, including rendered image, reward curve and success rate curve. The final success rates of all tasks are approaching 100% approximately.
IV-B Learning to Navigate in Cluttered Environment
The quadrotor should reach the target position safely while avoiding the obstacles during flight. It is one of the most fundamental tasks for quadrotors in real-world applications. We set a garage scene filled with random stones of density of $0.15/m^{3}$ in the spatial space as shown in Figure 5. It is more challenging more than dense scenes of trees or columns, since the quadrotor considers the vertial motion as well. Quadrotor starts from initial position, flies across the stone region and finally pass the target position. It takes $64\times 64$ depth image as visual observation and final position as target observation. This training takes 0.73 hour. To the best of our known, this instance is possible to be the first end2end autonomous navigation in cluttered environment without any training tricks, curriculum, privileged information or customized primitives within simulation. It is noted that, not the elegant reward setting but the abundant data contributes to the training.
IV-C Cooperatively Learning to Cross the Gap
We inserted two walls in the clear garage to create a narrow gap between them as Figure 5. Three quadrotors start from one side of the wall, fly towards distributed targets in the other side, and cross the narrow gap simultaneously. They must cooperate by considering each other’s state, to safely avoid both obstacle and other agents. Quadrotors take $64\times 64$ depth image as visual observation, final position as target observation, additionally other agents’ state as swarm observation. Unlike single-agent tasks, VisFly allows only one set of swarm in one scene, which limits the maximum batch size due to memory constraints. This training takes 1.1 hour. The result shows that, the by observing the other agents’ state, quadrotors are able to avoid collision with obstacles as well as each other.
V Discussion
VisFly’s torch-based dynamics achieve differentiable physics at the expense of lower accuracy compared to professional simulators like Gazebo. However, this trade-off does not negatively impact Sim2Real, even in extreme scenarios such as quadrotor racing[27], where the quadrotor must fly close to its dynamic upper bound. Additionally, VisFly does not integrate high-level controllers such as model predictive controller (MPC) for two primary reasons. Firstly, these advanced controllers consume excessive computational resources in both simulation and real flight. Secondly, end2end learning is preferred as it more closely imitates natural decision-making processes. In future work, we will incorporate real-world dynamics into VisFly to make HITL available.
VI Conclusion
In this paper, we introduce a both photo-realistic and sampling-efficient quadrotor simulator VisFly. VisFly integrates differentiable dynamics with habitat-sim, solving the shortage of visual training data especially for reinforcement learning. It wraps whole environments as gym interfaces, making it available to directly implement the latest reinforcement learning algorithms, such as stable-baselines3. It offers parallelized simulation to perform multi-agent system as well as multi-scene system. Since habitat-sim is widely used, it can also leverage a large amount of open-source scene datasets specifically designed for embodied learning, without pre-processing. We conducted three learning tasks involving visual input: learning to land, learning to navigate through a cluttered environment, and swarm learning to cross a narrow gap. The results confirm the reliability of this framework and provide foundational instances that enable users to extend high-level tasks. According to our current understanding, it could be the first quadrotor simulator that intensively generates available visual data in a short time. It is hopeful that VisFly could not only bridge the gap between simulation and reality but also between embodied learning and aerial robotics.
Acknowledgments
This work is inspired by Habitat-sim[20]. Part of the code architecture refers to Flightmare[17]. We sincerely appreciate their contributions to the quadrotor and embodied learning communities.