### v3

# VisFly: An Efficient and Versatile Simulator for Training Vision-based Flight

Abstract
We present VisFly, a quadrotor simulator designed to efficiently train vision-based flight policies using reinforcement learning algorithms. VisFly offers a user-friendly framework and interfaces, leveraging Habitat-Sim’s rendering engines to achieve frame rates exceeding 10,000 frames per second for rendering motion and sensor data. The simulator incorporates differentiable physics and seamlessly integrates with the Gym environment, facilitating the straightforward implementation of various learning algorithms. It supports the direct import of all open-source scene datasets compatible with Habitat-Sim, enabling training on diverse real-world environments and ensuring fair comparisons of learned flight policies. We also propose a general policy architecture for three typical flight tasks relying on visual observations, which have been validated in our simulator using reinforcement learning. The simulator will be available at [https://github.com/SJTU-ViSYS/VisFly].
Index Terms:
Quadrotor Simulation, Vision-based flight, learning-based navigation.
I Introduction
Quadrotors have great potential applications in various tasks such as search-and-rescue, surveillance, and inspection operations. The capability of autonomous flight of quadrotors is the key to enabling them to accomplish complex tasks, especially in challenging environments where GPS signals are unavailable and full of obstacles. Typically, an autonomous flight stack consists of localization, environmental perception, planning, and control components. Significant progress has been made in each component that allows quadrotors being able to autonomously navigate in cluttered environments [1, 2] with only onboard sensory and computational resources. However, the classic pipeline heavily relies on the performance of individual components. The accumulation of errors and running time of individual components could greatly reduce the agility and robustness of autonomous flight [3].
Table I: Key Character Comparison of Popular Simulators
Engine Parallel FPS Sensor Gym- 3D Different- Swarm
Physics Render agent scene Physics Render wrap Datasets iable
RotorS[4] Gazebo OpenGL $1\times 10^{2}$ 20 I, D, C
Airsim[5] PhysX Unreal $1\times 10^{2}$ 60 I, D, C, S
FlightGoggles[6] Flexible Unity $1\times 10^{3}$ 64 I, D, C, S
FastSim [7] Flexible Unity $1\times 10^{2}$ 40 I, D, C, S
CrazyS [8] Gazebo OpenGL $1\times 10^{3}$ 20 I, D, C
Flightmare[9] Flexible Unity $2.2\times 10^{5}$ $2.3\times 10^{2}$ I, D, C, S
PyBulletDrone[10] Bullet OpenGL $1\times 10^{4}$ $3\times 10^{3}$ I, D, C, S
OmniDrone[11] Self-defined Isaac Sim $2\times 10^{5}$ $3\times 10^{2}$ I, D, C, S, L
Ours Self-defined Habitat-sim $3.7\times 10^{4}$ $1\times 10^{4}$ I, D, C, S
  • 1
    I, C, D, S, and L refer to IMU, RGB, depth, segmentation, and LiDAR, respectively. Simulators marked with an asterisk (*) are full-stack, while the others are learning-specialized. VisFly is lightweight and does not have high hardware requirements. Note that the FPS data provided by the original papers were not evaluated with the same number of agents, but they still offer a performance reference. A dash (”-”) in the FPS column indicates that this function is unavailable. All frame rates are tested with 100 environments if parallelization is supported, and render FPS are tested while rendering $64\times 64$ depth images. FastSim cannot be deployed locally at present, so the FPS data is directly cited from the original paper. All frame rates are tested on a desktop with an RTX 4090 and a 32-core 13th Gen Intel(R) Core(TM) i9-13900K.
An alternative approach is to train a neural network (policy) to receive the sensory input and directly send the control commands to the quadrotor. Such methods, also known as the learning-based approach, typically integrate several modules into a single flight model, offering the potential to achieve autonomy with low latency and high robustness. One way to train the flight policy is to collect real-world data by human-pilot drones [12, 13, 14], but this approach is inefficient and prohibitively expensive. Alternatively, simulation can generate datasets, and policies trained on these simulated datasets often achieve performance comparable to those trained on real-world data[15]. Researchers [3] have demonstrated impressive collision avoidance capabilities by employing an algorithmic expert to gather training data for the flight policy. However, designing such experts to achieve both algorithmic optimality and efficiency can be challenging, thereby limiting the applicability of this approach to complex tasks.
Recently, researchers have demonstrated that reinforcement learning (RL) facilitates not only high-level decision-making applications such as autonomous driving [16], strategy games[17], but also basic control problems for complicated systems, like humanoid [18], quadruped locomotion [19]. These potential and powerful applications inspire academics in the field of aerial robotics to explore reinforcement learning to enhance autonomous flight [20, 21]. The RL-based approach usually relies on a simulator to collect massive data generated from the interaction between the agent and the simulated environment. There are already well-designed simulators [22, 5] widely used in the robotics community, with some specifically dedicated to quadrotors [4, 8]. These simulators are primarily designed for fast prototyping and iteration, typically encompassing the full stack of hardware and software simulations with an emphasis on closed-loop validation and testing. They usually operate at slow frame rates (100 fps), which does not negatively impact testing or validation but presents significant drawbacks for reinforcement learning due to low sampling efficiency. Some quadrotor simulators [9, 11, 10] have addressed these drawbacks by improving the sampling of physical state data, enabling the training of state-based policies for autonomous flight. However, they still lack the capability to sample vision input at high frame rates, making it challenging to train a high-performance vision-based flight policy.
Given that vision is one of the most informative and cost-effective sources for perception, it holds substantial potential for deploying a vision-based autonomous flight model in real-world applications. The key to effectively training a vision-based flight policy for quadrotors via reinforcement learning lies in 1) the rendering speed of the vision data, 2) the efficient computation of flight physics, 3) the diversity of scenes for training, and 4) the adoption of state-of-the-art reinforcement learning algorithms. Moreover, to enable the autonomous flight of a team of quadrotors, multi-agent learning should also be supported.
To achieve these goals, we developed an efficient simulation environment for training vision-based autonomous flight of quadrotors. Our environment leverages the PyTorch, Habitat-Sim [23], and OpenAI’s Gym [24] platforms. We implemented flight dynamics and three widely-used flight controllers using PyTorch, making flight physics not only computationally efficient but also differentiable, thereby enabling the use of model-based algorithms with known dynamics [25]. By adopting Habitat-Sim’s rendering engine, our environment can achieve high frame rates at 10,000 frames per second while rendering color images at a resolution of 64×64, which is about two orders of magnitude faster than existing traditional quadrotor simulators. Moreover, our simulator can load a series of open-source datasets comprising a vast number of 3D scene models (including real-world 3D scans), allowing us to train vision-based flight policies in diverse and realistic scenes. Our simulator is highly flexible, featuring gym-wrapped interfaces for integrating the latest learning algorithms and customized tasks. With our well-designed parallel mechanism, we can run multiple agents within different scenes simultaneously, facilitating training vision-based flight policies in multi-agent settings.
To validate our environment, we present an example of using a general network architecture to train vision-based flight policies for three autonomous flight tasks: landing, navigation, and cooperative flight through narrow gaps. All tasks rely solely on self-provided information, including vision and sensory inputs, without any external guidance. The results demonstrate that our simulation environment can be used to train vision-based flight policies in both single and multi-agent settings with high success rates in a very short period. Our simulation environment will be released as an open-source project and will be continuously updated. To highlight the novel features of our simulator, we compare it with existing simulators in Table I. To summarize, the contributions of this work include:
  1. 1.
    We present a novel quadrotor simulator that achieves high performance in rendering both physics and vision data, allowing to train vision-based flight policies efficiently.
  2. 2.
    Our simulator supports single-agent and swarm scenarios in parallelized physics and scenes, with Gym-wrapped interfaces for implementing the latest learning algorithms and customization of tasks. It also integrates open-source datasets for diverse scenes.
  3. 3.
    We demonstrate the effectiveness of our training environment for training diverse vision-based tasks, highlighting its versatility.
II Related Work
II-A Full-stack simulators
We compare the key factors of primary simulators in Table I, which provides a quick overview of the progress. The initial aim of simulators is to create a simulated platform for users to verify the algorithmic prototypes because quadrotors with immature algorithms are easy to crash. This category always consists of comprehensive full-stack simulated flight modules such as perception, planning, and control, especially for highly realistic physics. Gazebo [22] is a widely used full-stack robotics simulator of this category. While it is a general-purpose simulator, it has strong support for UAVs through various plugins and models. Considering various requirements, researchers develop simulators specific for UAVs based on Gazebo. RotorS [4] is a popular multi-rotor UAV modular simulator built on Gazebo, providing flexible interfaces to replace several parts of the pipeline, making it convenient to verify state estimators or control algorithms. CrazyS [8] is a lightweight simulator specialized for a software-to-hardware solution Crazyflie 2.0, which helps to easily deploy in the real world. Rather than using open-source rendering engines as in Gazebo, several simulators utilize commercial rendering engines such as Unity  [^1] https://unity.com/ and Unreal Engine  [^2] https://www.unrealengine.com/ to achieve high visual fidelity. For instance, FlightGoogles[6] employs the Unity engine for realistically simulating image sensors and uses real-world flight data to create avatars in simulation environments, enabling vehicle-in-the-loop simulation. Similarly, AirSim [5] is built on Unreal Engine to offer physically and visually realistic environments, along with real-time hardware-in-the-loop (HITL) simulation capabilities. Additionally, AvoidBench [26], built on the Unity engine, offers valuable benchmarks to evaluate obstacle avoidance algorithms. FastSim [7], also built on Unity, stands out as the first to integrate path planning algorithms directly into the simulator, facilitating tests of high-level autonomous control functionalities. While using commercial engines significantly enhances image quality (including depth maps and semantic segmentation), it comes at the cost of slower rendering speed because of their sophisticated rendering pipelines, particularly for the vision data. However, for applications focusing on learning in simulators using vision sensors, it is critical to acquire vision data quickly for fast training.
II-B Learning-specialized simulators
High acquisition rates for both sensory data and robot physics significantly accelerate the training process of learning-based algorithms. This makes them essential features for simulators designed for learning. Notably, physics computation can be achieved at very high frequencies in some existing simulators. For instance, by paralleled running, the frame rate of physics computation can reach an impressive 100000 FPS in Flightmare [9] and OmniDrone [11]. However, the frame rates will drop down to hundreds of FPS if the visual observation is rendered simultaneously in Flightmare. PyBulletDrone [10] achieves high rendering FPS via simple rendering engine OpenGL3 and TinyRender  [^3] https://github.com/ssloy/tinyrenderer. However, the direct basic OpenGL-based rendering engine lacks the functionality to import complex scenes and objects, making it difficult to train navigation policies using diverse and complicated scene data.
Selecting a rendering platform for training vision-based flight policies necessitates balancing speed, open-source scene availability, and hardware requirements. Habitat-sim[23], a lightweight simulator for training embodied agents, excels in rendering speed due to its customized rendering engine. It is compatible with numerous popular open-source datasets, including photo-realistic 3D reconstructions and digital forms, and offers flexible interfaces for customizing scenes and agent configurations, all with minimal hardware performance requirements. However, it focuses solely on ground robots and manipulations. In contrast, Nvidia’s Isaac Sim  [^4] https://developer.nvidia.com/isaac/sim is a comprehensive simulator that features extensive joint optimizations in rendering and simulation at the software architecture level. Nevertheless, it has higher hardware requirements and provides less flexibility and extensibility compared to Habitat-sim. Given these considerations, Habitat-sim appears to be a promising tool for addressing the limitations of existing learning-specialized simulators for aerial robots.
Our simulator is built on Habitat-sim, utilizing its fast rendering engine and well-designed interfaces to leverage existing massive 3D datasets. However, unlike Habitat-sim, which focuses on ground robots and manipulations, our simulator is tailored specifically for aerial robots. For the convenience of implementation, existing learning-specialized simulators, such as Flightmare, OmniDrone, and PyBulletDrone wrap the environment upon OpenAI’s gym environment standards. An OpenAI’s gym class receives an action, updates its internal state, and returns a new state paired with a corresponding reward. Like most of existing simulators, VisFly provides examples and tutorials to help beginners get started quickly. However, unlike others that primarily focus on state-only tasks—using only the UAV’s state, such as orientation, position, and velocities for control or planning—VisFly’s examples are specifically designed for vision-based flights. While some simulators, like OmniDrone, include examples using LiDAR inputs, VisFly emphasizes using visual data. Our examples are released with a modular design and are seamlessly integrated with Gym environments, providing users with an easy way to customize their vision-based flight tasks.
Figure 1: Overall diagram of VisFly. Our system uses differentiable physics to drive single or multiple quadrotor agents through four different types of controllers and employs Habitat-sim’s rendering engine for fast rendering and access to open-source datasets. All of these components are integrated into Gym environments, providing standard interfaces for various learning algorithms.
III Methodology
Here, we introduce three key features of VisFly and how they are implemented: (1) differentiable, and parallelized dynamics computations; (2) a rapid-rendering, parallelized, photo-realistic scene manager; (3) Gym-wrapped interfaces and extensions based on modular designs. A schematic diagram of VisFly is shown in Figure 1.
III-A Differentiable dynamics based on PyTorch
The dynamics simulation is implemented using PyTorch to enable efficient, parallel physics computations on GPUs. PyTorch’s computational graph also enables the computation of analytical gradients through back-propagation, making the dynamics computation differentiable. The differentiable dynamic model can be directly integrated into model-based reinforcement learning frameworks[27]. By following the exact dynamic process, analytical gradients offer more precise descent directions, significantly accelerating training compared to common value-based or policy-based reinforcement learning methods[25, 28]. We model quadrotor’s dynamics as follows:
$\dot{\mathbf{x}}_{W}={\mathbf{v}}_{W}$ (1)
$\dot{\mathbf{v}}_{W}=\frac{1}{m}\mathbf{R}_{WB}(\mathbf{f}+\mathbf{d})+\mathbf{g}$
$\dot{\mathbf{q}}=\frac{1}{2}\mathbf{q}\otimes\mathbf{\Omega}$
$\dot{\mathbf{\Omega}}=\mathbf{J}^{-1}(\eta-\mathbf{\Omega}\times\mathbf{J}\mathbf{\Omega})$
where $\mathbf{x}_{W}$, $\mathbf{v}_{W}$, $\mathbf{q}$ represent the position, velocity, and orientation (quaternion) of quadrotors in the world frame, and $\mathbf{\Omega}$ denotes the angular velocity in the body frame. The quaternion-vector product is represented by $\otimes$. The rotation matrix from the body frame to the world frame is $\mathbf{R}_{WB}$. The collective rotor thrust and the momentum are denoted by $\mathbf{f}$ and $\mathbf{\eta}$, respectively. The diagonal inertia matrix is $\mathbf{J}$, and the quadrotor mass is $m$. The air drag in the body coordinate, assumed proportional to the square of velocity [29], is represented by $\mathbf{d}$ :
$\small\mathbf{d}=\frac{1}{2}\rho\mathbf{v}_{B}\odot\mathbf{v}_{B}\mathbf{C}_{d}\odot\mathbf{s}$ (2)
where $\rho$, $\mathbf{C}_{d}$, $\mathbf{s}$, $\mathbf{v}_{B}$ denote the air density, drag coefficients, crossing area, and velocity in the body frame, respectively, the element-wise multiplication is denoted by $\odot$. The collective thrust and momentum, $\mathbf{f}$ and $\mathbf{\eta}$, are computed from individual rotor thrusts, $\mathbf{f}_{i}$:
$\small\mathbf{f}=\sum_{1}^{4}f_{i},\quad\mathbf{\eta}=\sum_{1}^{4}\mathbf{T}_{i}\times\mathbf{f}_{i}.$ (3)
Here $\mathbf{T}_{i}$ denotes the moment arm matrix for the rotor $i$. To account for control input delay, we introduce the time decay constant $c$ and model it as an exponential process:
$\small\mathbf{f}_{i}(3)=k_{2}\omega_{i}^{2}+k_{1}\omega_{i}+k_{0},\quad\omega_{i}=\omega_{i}^{des}+(\omega_{i}^{\prime}-\omega_{i}^{des})e^{-ct}$ (4)
where $\omega_{i}$ denotes the spinning speed for rotor $i$, and $\omega_{i}^{\prime}$, $\omega_{i}^{des}$ indicate the current and desired states, respectively, the constants $k_{2},k_{1},k_{0}$ are thrust coefficients with respect to rotor speeds. Based on [30, 31], we develop three widely-used control methods: individual single-rotor thrusts (SRT), mass-normalized collective thrust and body rates (CTBR), position commands and yaw (PS), and linear velocity commands and yaw (LV). SRT and CTBR are control inputs for Betaflight, which can be used for training efficient end-to-end policies[32]. The minimal simulation time step can be set several times shorter than the control time step, resulting in more precise physics. Both the fourth-order Runge-Kutta and Euler methods are supported, providing users the flexibility to choose based on their platform’s performance.
Figure 2: RGB, depth, and semantic segmentation captured in open-source and customized scenes. On the right is a swarm of quadrotors in a clear garage.
III-B Efficient rendering and open-source scene management
We use the rapid rendering engine of Habitat-sim to render the visual observation data efficiently. Figure 3 presents the frame rate of VisFly while running physics simulations. With visual input, VisFly can achieve a frame rate of up to 10,000 FPS, almost 100 times faster than simulators using Unity or Unreal Engine.
Figure 3: Frame rate performance of VisFly. Left: Tested with 100 agents in the Replica dataset, achieving up to 10,000 FPS. At 256$\times$256 resolution, the frame rate still reaches up to 6,000 FPS. Right: Variation in frame rate with the number of scenes running simultaneously at 64$\times$64 resolution.
Similar to Habitat-Sim, VisFly supports a range of open-source 3D scene datasets, including HM3D[33], MP3D[34], Replica[35], and HSSD[36], many featuring highly photorealistic real-world scans. These datasets facilitate direct training and evaluation of models. To augment these resources, we incorporate a scene generator for producing random, cluttered environments, such as those used in collision avoidance training. The combination of real-world data and customizable scenes enhances VisFly’s adaptability, reducing Sim-to-Real challenges. By eliminating the need for extensive data collection, these datasets also foster fair comparisons and establish strong baselines for quadrotor research. VisFly supports common vision and motion sensors, and is extensible to complex multi-camera systems.
One requirement in training flight policies is to query the nearby obstacles on the fly. Habitat-sim, specialized for ground robots, assumes that robots operate on a 2-dimensional surface. It precomputes the NavMesh, which projects spatial objects onto the floor and generates distance information to the closest obstacle. This approach is not suitable for aerial robots that move in 3D space. Therefore, we developed a spatial distance computation module via the CGAL [^5] https://www.cgal.org/ package in C++, enabling it to detect the relative position of the nearest point on the obstacle without collision actually happening.
VisFly supports parallelized agents as well as multi-agents within the same scene. These two modes share the same underlying framework. The key difference between them is that parallelized agents visually omit other drones and ignore collisions between them, while multi-agent scenarios include interactions and collisions among all drones.
III-C Modular design for Gym integration
Gym [24] is an open-source toolkit providing a standard interface for reinforcement learning. It offers a collection of pre-built environments and a framework for custom environment development. Agents are able to interact with environments through step() function, obtaining rewards, observations, termination signals, and other information at each time step.
VisFly’s ultimate mission is to facilitate the development of aerial robot learning for autonomous flight tasks. Therefore, we have taken the extensibility of the basic environment into account as much as possible when packaging the environment. Two abstract environment classes droneGymEnvsBase, multiDroneGymEnvsBase are provided. When users design custom tasks for a quadrotor, they need only define a class that inherits from the abstract gym environment class and overrides three functions: get_success(), get_reward(), and get_observation() as shown in Lst. 1. This approach creates a new gym environment tailored to the specific mission. We provide three examples in Section IV for users to reference. This hierarchical architecture enables the training of models that can act as backbones for more complex tasks. For instance, a model trained for obstacle avoidance can serve as the low-level policy or backbone for higher-level autonomous tasks, such as search and rescue operations in wilderness areas.
1class DroneGymEnvsBase:
2 @abstract method
3 def get_reward()
4 @abstract method
5 def get_observation()
6 @abstract method
7 def get_sucess()
8 ...
9class DroneNavigationEnvs(DroneGymEnvsBase)
10 def get_reward()
11 ...
12 return r_distance+r_speed-r_collision
13 def get_observation()
14 ...
15 return state+vision+target
16 def get_sucess()
17 ...
18 return status_reach_the_target
19 ...
20class DroneLandingEnvs(DroneGymEnvsBase)
21 def get_reward()
22 ...
23 return -r_height+r_speed-r_collision
24 def get_observation()
25 ...
26 return state+vision+visual_target
27 def get_sucess()
28 ...
29 return status_height<0.1&status_speed<0.1
Lst. 1: Mocular class design for different tasks
III-D Domain randomization for sensor data and flight states
The gap between simulation and reality usually hinders the real-world deployment of models trained in simulation. In order to reduce the sim-to-real domain gap, one effective approach is to improve the realism of simulation, whose drawback is discussed in Introduction. The other approach is to enlarge the data distribution generated by simulators, strengthening the in-distribution generalization of models, which is called domain randomization[15]. For training vision-based flight policies, it is important to add noises to both the motion and vision sensors as listed in Table II.
Table II: Noise model of sensors
Sensor Noise Model
IMU Normal
RGB Normal, Poisson, Salt&Pepper, Speckle
Depth Normal, Poisson, Salt&Pepper, Speckle, RedWood[37]
Segmentation Normal, Poisson, Salt&Pepper, Speckle
In addition to adding noises to sensors, VisFly allows for the randomization of initial conditions, including position, velocity, orientation, and angular velocity, which can follow either a normal or uniform distribution. To achieve better generalization of the policy, agents are reset with initial positions away from obstacles in cluttered environments. An important contribution of VisFly is its capability to run multiple scenes distributively and aggregate observations centrally. During training, scenes are treated as datasets, allowing them to be sampled and shuffled. To the best of our knowledge, VisFly is the first simulator that supports simultaneous execution of multiple scenes for enhanced scene randomization.
IV Training examples
To validate this simulator, we present three cases of training vision-based flight policies: (1) learning to land, (2) learning to navigate in a cluttered environment, and (3) learning to cross the narrow gap cooperatively. We use reinforcement learning to train end-to-end models for controlling the drones using the vision inputs. To focus on training efficiency rather than the algorithms themselves, we employ PPO [38] from stable-baselines3[39] (SB3) for its outstanding capability to handle high-dimensional observationsThe backbone architecture of the policy network follows the standard design in SB3 , as shown in Figure 4, with modifications to enhance its adaptability across various tasks.
The image feature extractor is a conventional neural network that comprises three layers. Each layer includes a convolutional layer, an activation function, and a max pooling operation, followed by a linear layer with 128 units. The convolutional layers have kernel sizes of 5, 3, and 3, with channel counts of 6, 12, and 18. Each max pooling operation uses a kernel size of 2. The target extractor, state extractor, and swarm extractor are all multilayer perceptrons (MLPs) with two linear layers, having 128 and 64 units. The action net and value net are also MLPs, with two linear layers of 64 units. All activation functions are ReLU. This architecture has been parameterized, making it convenient to fine-tune. The action type in all the examples is CTBR. The quadrotor takes its full state (position, orientation, linear velocity, angular velocity), visual image, and target as observations. Most of the training hyper-parameters are set to default values. Detailed settings, the reward, done, and observation functions are provided in the source code.
Figure 4: Neural network architecture used across different tasks. The image feature extractor has been modified to be more customizable, including backbones such as ResNet[40], MobileNet[41], EfficientNet[42]. The policies for the three tasks follow this architecture but are slightly different. VisFly additionally incorporates a recurrent network interface [43] for training tasks that do not strictly conform to the Markov process.
IV-A Learning to land
In this task, the quadrotor starts in flight, gradually decelerates, and eventually lands safely on the landing area. It continuously monitors the limited landing area directly below, which is represented as a black rectangle positioned horizontally on the ground within a clear garage, as shown in Figure 5. The quadrotor calculates the center of the $0.5\times 0.5m^{2}$ black rectangle in its field of view, strives to keep it centered, and incorporates it into its observation. During the flight, the quadrotor captures only $64\times 64$ RGB image from the camera rather than acquiring precise spatial information about the landing area. The center of the landing area is determined using threshold segmentation and included as a target observation. This training process takes 1.7 hours, with additional time required for prepossessing the center of the landing area, as detailed in Section IV-B.
Figure 5: Training results of landing, navigation, and swarm crossing. These subfigure columns respectively represent the results of landing, navigation, swarm crossing, including rendered image, reward curve and success rate curve. The final success rates of all tasks are approaching 100% approximately.
IV-B Learning to navigate in a cluttered environment
In this task, the quadrotor must reach the target position safely while avoiding obstacles during flight. It is one of the most fundamental tasks for quadrotors in real-world applications. We set a garage scene filled with random stones of the density of $0.15/m^{3}$ in the spatial space as shown in Figure 5. This scenario is more challenging than dense scenes with trees or columns because it also requires the quadrotor to account for vertical motion. The quadrotor starts from an initial position, navigates through the stone-filled region, and eventually reaches the target position. It uses a $64\times 64$ depth image as the visual observation and the final position as the target observation. Training takes 0.73 hours. To our knowledge, this is the first instance of end-to-end autonomous navigation in a cluttered environment without employing training techniques such as curriculum learning, privileged information, or customized primitives within the simulation.
IV-C Learning to cross a narrow gap cooperatively
We inserted two walls in the clear garage to create a narrow gap between them as Figure 5. Three quadrotors start from one side of the walls, fly towards distributed targets on the other side, and pass through the narrow gap simultaneously. They must cooperate by considering each other’s state to avoid obstacles and other agents safely. Each quadrotor uses a $64\times 64$ depth image as its visual observation, the final position as its target observation, and the states of other agents as its swarm observation. Unlike single-agent tasks, VisFly supports only one swarm per scene, which limits the maximum batch size due to memory constraints. Training takes 1.1 hours. The results show that by observing the states of other agents and the environment, the quadrotors can effectively fly across the gap and avoid collisions with each other.
V Discussion
VisFly’s PyTorch-based dynamics provide differentiable physics, though with slightly lower accuracy compared to professional simulators like Gazebo. However, lower accuracy does not impair the Sim-to-Real deployment. This has been demonstrated in challenging scenarios such as quadrotor racing [20] where the quadrotor is capable of flying close to its dynamic limits. Additionally, VisFly does not incorporate high-level controllers like Model Predictive Control (MPC) for two main reasons: First, these advanced controllers require substantial computational resources during simulation and real flight. Second, End-to-end learning is preferred because it closely mimics natural decision-making processes. In future work, we plan to integrate real-world dynamics into VisFly to enable Hardware-in-the-Loop (HITL) capabilities.
VI Conclusion
This paper presents VisFly, a quadrotor simulator designed for efficient training of vision-based flight policies. Leveraging differentiable physics and Habitat-Sim’s rendering capabilities, VisFly supports single and multi-agent simulations across various controller types. Integrating with the gym environment provides a standardized interface for various reinforcement learning algorithms. Experimental results on landing, cluttered environment navigation, and cooperative gap crossing demonstrate the simulator’s effectiveness. VisFly serves as a robust foundation for future research, offering baseline implementations and extensibility for advanced quadrotor tasks.